# Spin-free orbital entropy, mutual information, and entanglement analysis

### v1

@@ Abstract
Orbital entropies, pair entropies, and mutual information have become popular tools for analysis of strongly correlated wave functions. They can quantitatively measure how strongly an orbital (e.g. from the DMRG active space) participates in the strong correlation and reveal the entanglement pattern between different orbitals. However, this pattern can become rather complicated and sometimes difficult to interpret for large active spaces and is not invariant with respect to the spin projection ($M_{s}$) component of the spin multiplet state. We introduce a modified spin-free orbital entropy, pair entropy, and mutual information, which simplify the entanglement analysis and are invariant with respect to $M_{s}$. By comparison of these quantities with their “original” spin-including counterparts one can distinguish static correlation due to spin couplings from the “genuine” strong correlation due to a multiconfigurational character of the wave function. We illustrate the approach on a model consisting of a non-interacting dimer of triplet diradicals and on a more realistic example of iron-sulfur bound complexes with one and two iron atoms.
@@ 1 Introduction
The problem of electron correlation has a central importance in quantum chemistry. The correlation effects are often divided into two or three categories, although the division is somewhat fuzzy and not completely rigorous. The dynamic (weak) correlation occurs due to pairwise Coulomb repulsion of the electrons when they come close to each other and corresponds to many small contributions of excited Slater determinants to the wave function. On the other hand, one speaks about strongly (or non-dynamically) correlated systems, when there is no single dominant Slater determinant, and two or more determinants have a comparable weight. Sometimes one distinguishes two sub-cases — static correlation, where several Slater determinants with large coefficients are needed to obtain correct spin symmetry, and “true” strong correlation which is related to quasidegenerate orbitals in molecules undergoing bond breaking, in transition metal complexes, etc. For systems exhibiting dynamic (weak) correlation only, many single-reference methods exist, which are able to achieve a satisfactory “chemical accuracy”. For strongly correlated systems, the situation is much more difficult, since one has to accurately take into account both the strong and the dynamic correlation. Traditionally, the presence of the strong correlation has been judged based on the magnitude of configuration interaction (CI) coefficients or the size of coupled cluster (CC) amplitudes. New tools for this purpose have been developed based on the notion of entanglement, which is central to quantum information theory [1, 2, 3]. The analysis employs the orbital entropy, which is a measure of the entanglement of a given orbital with all other orbitals, and the mutual information, which measures the entanglement of two orbitals with each other, regardless of other orbitals. These approaches have become more popular in recent two decades [4, 5, 6, 7, 8, 9, 10], since they are particularly suitable for use with the DMRG method, which also gained popularity in quantum chemistry, particularly for studies of strongly correlated transition metal compounds [11, 12, 13, 14, 15, 16]. Iron-sulfur clusters and complexes represent an important class of such compounds which is highly relevant in biochemistry [15, 17, 18]. They facilitate redox chemistry in living matter by carrying out electron transfer and catalytic processes essential e.g. for nitrogen fixation, respiration, and photosynthesis. Their computational treatment represents a challenge due to the presence of strong correlation and many close lying states of different spin multiplicity resulting from the partially occupied 3d-shells of iron atoms. Although the high-spin state in these systems is dominated by a single configuration (but its lower-$M_{s}$ components are still multideterminantal), the dynamic correlation can energetically favor the low-spin states, which typically have more complicated multiconfigurational character[16, 19]. In this work we introduce spin-free orbital entropy and mutual information and illustrate their application to the entanglement analysis in these systems.
@@ 2 Theory
Orbital entanglement measures have been recognized as an important tool for analysis of correlated electronic wave functions already in the nineties [1, 2, 3] and have since become widely used [4, 7, 8, 9, 10], hand-in-hand with the development of the DMRG method in quantum chemistry [20].
The central quantity in the quantum information theory is the von Neumann entropy defined as
@@
Due to the sub-additivity of the von Neumann entropy $I_{AB}\geq 0$ [21]. For pure states, the equality is achieved only if the state of AB is separable and the mutual information can thus provide a measure of entanglement in this context. However, it can be nonzero for separable mixed states, reflecting their classical correlation.
@@
The sum of all orbital entropies (in the active space of $k$ orbitals) is often called total entropy
@@
where $\omega_{p}$ are the eigenvalues of $\rho_{ij}$ (we omit the indices $i,j$ at $\omega_{p}$ for simplicity). The mutual information between orbitals $i$ and $j$ is then in quantum chemistry usually defined as
$I_{ij}=\frac{1}{2}(S_{i}+S_{j}-S_{ij})(1-\delta_{ij})\;\;.$ (11)
This differs from the definition (3) by a factor and by the Kronecker $\delta$ term assuring that $I_{ii}=0$. Using these quantities, a weighted complete graph on $k$ vertices can be constructed, $S_{i}$ being vertex weights and $I_{ij}$ serving as edge weights. A picture of this graph with weight coded by color and/or line thickness then illustrates the entanglement in the molecule (for a particular choice of orbitals). As a measure of the overall entanglement in a given state, the sum of pairwise mutual informations can be employed
@@
The reduced density matrices can then be efficiently computed by contractions of the density matrix $|\Psi\rangle\langle\Psi|$ expressed using MPS [23]. The entanglement analysis thus became more popular in quantum chemistry hand in hand with the advent of DMRG.
Transition metal complexes, and particularly the multi-metallic ones, are examples of molecules with the most complicated strongly correlated electronic structure. Even for a single electron configuration (a given occupation of the open shells) there are in general are many spin-adapted configurations state functions (CSF) and even more Slater determinants. The entanglement analysis as described above, is sensitive to the electron spin couplings, as the ”up” and ”down” singly occupied orbital represent two different basis functions in ${\cal B}$. This may lead to a rather complicated entanglement diagram, which cannot distinguish which part of the correlation is “static” due to the spin coupling and which is a “genuine” strong correlation. In the following we present a modified orbital entropy, pair entropy and mutual information, which are spin-free, i.e. consider a singly occupied orbital regardless of the direction of the spin of its electron. We start with the orbital basis (5) and consider the two singly occupied orbital states as a single indistinguishable “microstate”
@@
where the summation runs over the spin-free basis states from $\tilde{\cal B}\otimes\tilde{\cal B}$ and employs the “spin-summed eigenvalues” of $\rho_{ij}$
@@
Here $\omega_{p}$ are the “original” eigenvalues of $\rho_{ij}$ and $c_{pk}$ are the corresponding normalized eigenvectors, while ${\cal B}_{q}$ denotes the subset of the pair basis $\cal B\otimes\cal B$ which corresponds to the $q$-th element of $\tilde{\cal B}\otimes\tilde{\cal B}$. The quantity $\tilde{\omega}_{q}$ thus represents a sum of $\rho_{ij}$ eigenvalues weighted by the contribution of their eigenvector to a given spin-free basis state. Clearly, $0\leq\tilde{\omega}_{q}\leq 1$ and
@@
The $\tilde{\omega}_{q}$ quantities can thus be interpreted as probabilities and employed in the entropy expression (19). By the same argument as for orbital entropy, the spin-free pair entropy is not bigger than the original one $\tilde{S}_{ij}\leq S_{ij}$, with equality if all electrons have the same spin. The “spin-summed eigenvalues” (20) might seem a bit ad-hoc construction, but when reconstructing $\rho_{ij}$ from its eigenvalues and eigenvectors as
@@
since this trace combined with the trace in (23) forms a complete trace over orbital $j$ combined with the spin summation over orbital $i$. When the two orbitals $i$ and $j$ are totally uncorrelated, $\rho_{ij}=\rho_{i}\otimes\rho_{j}$ and due to the diagonality of the orbital RDMs, $\rho_{ij}$ is also diagonal with eigenvalues being products of the $\rho_{i}$ and $\rho_{j}$ diagonal elements, while the eigenvectors $c_{pk}$ form a unit matrix. From (20) then follows that the $\tilde{\omega}_{q}$ quantities are products of the spin-summed orbital RDMs
@@
$\tilde{I}_{ij}=\frac{1}{2}(\tilde{S}_{i}+\tilde{S}_{j}-\tilde{S}_{ij})(1-\delta_{ij})\;\;.$ (29)
The spin-free mutual information is never bigger than the original one $\tilde{I}_{ij}\leq I_{ij}$. This is a consequence of the data processing inequality from the classical information theory (cf. Theorem 2.8.1 in [24]). It says that if $x,y,z$ are probability distributions and $z=f(y)$, then for the mutual information holds $I(x,y)\geq I(x,z)$. Since the mutual information is symmetric, it can be applied twice and for $x^{\prime}=f(x)$ and $y^{\prime}=g(y)$ we have $I(x,y)\geq I(x^{\prime},y^{\prime})$. The orbital and pair entropies and consequently the mutual information (11,29) are computed from the eigenvalues of the density matrices, which fulfill all requirements on being interpreted as probability distributions, so the classical information theory results hold for these quantities. Since the spin summation (20) is a function mapping between two such distributions, the data processing inequality must hold. Intuitively, it is clear that if we discard some information by performing the spin summation, we cannot increase the amount of information correlating one orbital with another. The transition from the original mutual information (11) to the spin-free one (29) can thus only simplify the diagrams for the entanglement analysis.
Another advantage of the spin-free quantities is their invariance with respect to the spin component $M_{s}$ of a given spin multiplet state. To show this, we start from the well known fact that the total $N$-electron wave function can be written as a product of spatial and spin wave functions
@@
where $\Psi^{\prime}_{{\cal E}_{i}}(N-1)$ and $\Psi^{\prime\prime}_{{\cal E}_{i}}(N-1)$ are not in general spin eigenfunctions, the spin-trace over ${\cal E}_{i}$ does not form a full spin trace and the $(\rho_{i})_{22}$ and $(\rho_{i})_{33}$ matrix elements are not invariant. However, when taking the sum of the matrix elements $(\rho_{i})_{22}+(\rho_{i})_{33}$, we sum over the spin variable of the electron in orbital $i$ as well, forming a full trace over all the spin variables
@@
which manifestly expresses the invariance of this quantity (and consequently of the spin-free entropy $\tilde{S}_{i}$) with respect to $M_{s}$. Actually, this equation might serve as an alternative definition of the spin-free orbital RDM, which would be particularly suitable in the context of spin-adapted methods like GUGA-CI. For the pair reduced density matrix, the $M_{s}$ invariance can be shown analogously, analyzing the 16 cases corresponding to the pair basis ${\cal B}\otimes{\cal B}$. Notice, however, that neither the spin-free nor the spin-dependent orbital entropies and mutual informations are invariant with respect to orbital transformations, so the entanglement pattern depends on the form of the orbitals (e.g. localization) [25] and a sensible (chemically intuitive) choice can be very important for its interpretation [26]. Moreover, by a suitable orbital transformations (called fermionic mode optimization in DMRG [27]), it is possible to minimize the multireference character (i.e. the orbital entanglement) in the wave function [28].
@@ 3 Computational details
For the illustration of the spin-free entanglement analysis we employed a model example of a non-interacting dimer of CH2 diradicals in triplet ground states coupled to different resulting multiplicities. We took arbitrarily an equilateral triangular geometry with bond lengths 1.121Å  and angle 152.7 and computed FCI for the monomer (triplet ground state) employing canonical RHF orbitals in the STO-3G basis and freezing the 1s orbitals on carbon atoms. The dimer was constructed from two monomers separated by 100 a.u. distance, and through different choices of initial guess we converged the FCI to various linear combinations of the degenerate ground states with different spin multiplicities. We employed our in-house FCI and entropy analysis C++ code for these calculations.
@@
As a more realistic application we have chosen the iron complexes [Fe(SCH3)4]- and [Fe2S2(SCH3)4]2- (cf. Fig. 1), which were recently employed by Neese et al. to demonstrate the new restricted open shell SCF code in Orca [19]. The geometries of the complexes were originally reported in Refs. [17, 18], but are available in the Supplementary Information of Ref. [19]. We employed the Orca program [29] and the Def2-TZVP basis set to perform a configuration-averaged open shell SCF calculations and integral dump, while the MOLMPS code[30] was used for the subsequent DMRG calculations, and the new C++ code for the entropy analysis. For the single-iron complex, the SCF was done configuration-averaged for 5 electrons in the 5 d-orbitals and DMRG(13,14) calculations with bond dimension 512 were performed for states with multiplicities 2,4,6. For the double-iron-sulfur complex, the configuration-averaged SCF was performed for the 10 electrons in iron d-orbitals and subsequently DMRG (14,14) and (22,21) calculations with bond dimension 2048 have been computed for the spin multiplicities 1,3,…,11. It turns out that in this complex the dynamic correlation is critical to get the correct energetic ordering of the states; indeed, in the (14,14) the high-spin state is incorrectly the lowest, while already in the (22,21) space the energy ordering was qualitatively correct (singlet ground state), in agreement with the post-DMRG and NEVPT2 calculations [16]. However, it was not possible to converge all excited states corresponding to low $M_{s}$ components of the high $S$ states.

### v2

# Spin-free orbital entropy, mutual information, and correlation analysis

@@ Abstract
Orbital entropies, pair entropies, and mutual information have become popular tools for analysis of strongly correlated wave functions. They can quantitatively measure how strongly an orbital (e.g. from the DMRG active space) participates in the strong correlation and reveal the correlation pattern between different orbitals. However, this pattern can become rather complicated and sometimes difficult to interpret for large active spaces and is not invariant with respect to the spin projection ($M_{s}$) component of the spin multiplet state. We introduce a modified spin-free orbital entropy, pair entropy, and mutual information, which simplify the correlation analysis and are invariant with respect to $M_{s}$. By comparison of these quantities with their “original” spin-including counterparts one can distinguish static correlation due to spin couplings from the “genuine” strong correlation due to a multiconfigurational character of the wave function. We illustrate the approach on a model consisting of a non-interacting dimer of triplet diradicals and on a more realistic example of iron-sulfur bound complexes with one and two iron atoms.
@@ 1 Introduction
The problem of electron correlation has a central importance in quantum chemistry. The correlation effects[1] are often divided into two or three categories, although the division is somewhat fuzzy and not completely rigorous. The dynamic (weak) correlation occurs due to pairwise Coulomb repulsion of the electrons when they come close to each other and corresponds to many small contributions of excited Slater determinants to the wave function. On the other hand, one speaks about strongly (or non-dynamically) correlated systems, when there is no single dominant Slater determinant, and two or more determinants have a comparable weight. Sometimes one distinguishes two sub-cases — static correlation, where several Slater determinants with large coefficients are needed to obtain correct spin symmetry, and “true” strong correlation which is related to quasidegenerate orbitals in molecules undergoing bond breaking, in transition metal complexes, etc. For systems exhibiting dynamic (weak) correlation only, many single-reference methods exist, which are able to achieve a satisfactory “chemical accuracy”. For strongly correlated systems, the situation is much more difficult, since one has to accurately take into account both the strong and the dynamic correlation. Traditionally, the presence of the strong correlation has been judged based on the magnitude of configuration interaction (CI) coefficients or the size of coupled cluster (CC) amplitudes. New tools for this purpose have been developed based on the notions of entanglement and correlation, which are central to quantum information theory (QIT) [2, 3, 4, 5, 6, 7, 8]. The antanglement and correlation analysis employs the orbital entropy, which is a measure of correlation of a given orbital with all other orbitals, and the mutual information, which measures the correlation of two orbitals with each other, regardless of other orbitals. In this article we do not distinguish between quantum and classical correlation of the orbitals, which are both captured by the aforementioned quantities. The QIT-based approaches have become more popular in recent two decades [9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21], since they are particularly suitable for use with the DMRG method, which also gained popularity in quantum chemistry, particularly for studies of strongly correlated transition metal compounds [22, 23, 24, 25, 26, 27, 28, 29]. Among the important applications of the entropy-based measures belongs the optimization of orbital ordering, which is essential to make the DMRG calculations efficient [9, 30] and the automated approach to active space selection [31], which has the potential to turn the application of multi-configurational methods into a black box approach. Further development in this direction represents the Quantum Information-Assisted Complete Active Space Optimization [32], where besides the active space selection also an orbital optimization minimizing the out-of-CAS orbital correlation is performed.
Iron-sulfur clusters and complexes represent an important class of such compounds which is highly relevant in biochemistry [26, 33, 34]. They facilitate redox chemistry in living matter by carrying out electron transfer and catalytic processes essential e.g. for nitrogen fixation, respiration, and photosynthesis. Their computational treatment represents a challenge due to the presence of strong correlation and many close lying states of different spin multiplicity resulting from the partially occupied 3d-shells of iron atoms. Although the high-spin state in these systems is dominated by a single configuration (but its lower-$M_{s}$ components are still multideterminantal), the dynamic correlation can energetically favor the low-spin states, which typically have more complicated multiconfigurational character and represent thus a challenge to computational methods [29, 35, 36, 37]. In this work we introduce spin-free orbital entropy and mutual information and illustrate their application to the correlation analysis in these systems.
@@ 2 Theory
Orbital correlation measures have been recognized as an important tool for analysis of correlated electronic wave functions already in the nineties [2, 3, 4] and have since become widely used [9, 13, 15, 16, 18], hand-in-hand with the development of the DMRG method in quantum chemistry [38]. They are, however, not restricted to any particular wave function ansatz and can be used with other strongly correlated methods as well, for example pCCD [19, 20, 21]. The central quantity in the quantum information theory is the von Neumann entropy defined as
@@
Due to the sub-additivity of the von Neumann entropy $I_{AB}\geq 0$ [39]. For pure states, the equality is achieved only if the state of AB is separable (which for pure states means a product state) and the mutual information can thus provide a measure of entanglement in this context. However, it can be nonzero for separable mixed states, reflecting their classical correlation. In general, the mutual information mesures correlation of both classical and quantum nature. Recently, additional quantum-information-theoretic measures have been introduced, which serve to distinguish the quantum and classical correlations [40, 7, 41].
@@
The sum of all orbital entropies (in the active space of $k$ orbitals) is often called total quantum information
@@
where $\omega_{p}$ are the eigenvalues of $\rho_{ij}$ (we omit the indices $i,j$ at $\omega_{p}$ for simplicity). The mutual information between orbitals $i$ and $j$ is then
$I_{ij}=S_{i}+S_{j}-S_{ij}\;\;.$ (11)
Some works in quantum chemistry included in the definition of orbital mutual information (11) a factor $1/2$ and a Kronecker $\delta$ term assuring that $I_{ii}=0$ [13], however, it is now customary to follow strictly the quantum information theoretic formula (3). Using these quantities, a weighted complete graph on $k$ vertices can be constructed, $S_{i}$ being vertex weights and $I_{ij}$ serving as edge weights. A picture of this graph with weight coded by color and/or line thickness then illustrates the correlation in the molecule (for a particular choice of orbitals). As a measure of the overall correlation in a given state, the sum of pairwise mutual informations can be employed
@@
The reduced density matrices can then be efficiently computed by contractions of the density matrix $|\Psi\rangle\langle\Psi|$ expressed using MPS [43]. The correlation analysis thus became more popular in quantum chemistry hand in hand with the advent of DMRG.
Transition metal complexes, and particularly the multi-metallic ones, are examples of molecules with the most complicated strongly correlated electronic structure. Even for a single electron configuration (a given occupation of the open shells) there are in general are many spin-adapted configurations state functions (CSF) and even more Slater determinants. The correlations analysis as described above, is sensitive to the electron spin couplings, as the ”up” and ”down” singly occupied orbital represent two different basis functions in ${\cal B}$. This may lead to a rather complicated correlation diagram, which cannot distinguish which part of the correlation is “static” due to the spin coupling and which is a “genuine” strong correlation. In the following we present a modified orbital entropy, pair entropy and mutual information, which are spin-free, i.e. consider a singly occupied orbital regardless of the direction of the spin of its electron. We start with the orbital basis (5) and consider the two singly occupied orbital states as a single indistinguishable “microstate”
@@
where the summation runs over the nine spin-free basis states ~$\tilde{\cal B}_{q}$ from $\tilde{\cal B}\otimes\tilde{\cal B}$, i.e. ~$\tilde{\cal B}_{q}\in\{|00\rangle,|01\rangle,|02\rangle,|10\rangle,\ldots,|22\rangle\}$, where $|00\rangle$ abbreviates $|0\rangle\otimes|0\rangle$ etc.. The quantities $\tilde{\omega}_{q}$ in (19) are the “spin-summed eigenvalues” of $\rho_{ij}$
@@
Here $\omega_{p}$ are the “original” eigenvalues of $\rho_{ij}$ and $c_{pk}$ are the corresponding normalized eigenvectors, while ${\cal B}_{q}$ denotes the subset of the pair basis $\cal B\otimes\cal B$ which corresponds to the $q$-th element of $\tilde{\cal B}\otimes\tilde{\cal B}$. For example, for ~$\tilde{\cal B}_{q}=|1\rangle\otimes|2\rangle$, ${\cal B}_{q}=\{|\uparrow\rangle\otimes|\uparrow\downarrow\rangle,|\downarrow\rangle\otimes\uparrow\downarrow\rangle\}$.
The quantity $\tilde{\omega}_{q}$ thus represents a sum of $\rho_{ij}$ eigenvalues weighted by the contribution of their eigenvector to a given spin-free basis state. Clearly, $0\leq\tilde{\omega}_{q}\leq 1$ and
@@
The $\tilde{\omega}_{q}$ quantities can thus be interpreted as probabilities and employed in the entropy expression (19). By the same argument as for orbital entropy, the spin-free pair entropy is not bigger than the original one $\tilde{S}_{ij}\leq S_{ij}$, with equality if all electrons have the same spin. The “spin-summed eigenvalues” (20) might seem a bit ad-hoc construction, but actually it can be viewed as a generalization of the single-orbital spin-free entropy case (17), where the orbital RDM is diagonal, and hence the set of eigenvectors forms a unit matrix. Moreover, when reconstructing $\rho_{ij}$ from its eigenvalues and eigenvectors as
@@
since this trace combined with the trace in (23) forms a complete trace over orbital $j$ combined with the spin summation over orbital $i$. Notice that other choices of the basis states for the spin-free two-orbital RDM, for example the 5 block space by total particle number and $M_{s}$, would not fulfill this requirement. When the two orbitals $i$ and $j$ are totally uncorrelated, $\rho_{ij}=\rho_{i}\otimes\rho_{j}$ and due to the diagonality of the orbital RDMs, $\rho_{ij}$ is also diagonal with eigenvalues being products of the $\rho_{i}$ and $\rho_{j}$ diagonal elements, while the eigenvectors $c_{pk}$ form a unit matrix. From (20) then follows that the $\tilde{\omega}_{q}$ quantities are products of the spin-summed orbital RDMs
@@
$\tilde{I}_{ij}=\tilde{S}_{i}+\tilde{S}_{j}-\tilde{S}_{ij}\;\;.$ (29)
The spin-free mutual information is never bigger than the original one $\tilde{I}_{ij}\leq I_{ij}$. This is a consequence of the data processing inequality from the classical information theory (cf. Theorem 2.8.1 in [44]). It says that if $x,y,z$ are probability distributions and $z=f(y)$, then for the mutual information holds $I(x,y)\geq I(x,z)$. Since the mutual information is symmetric, it can be applied twice and for $x^{\prime}=f(x)$ and $y^{\prime}=g(y)$ we have $I(x,y)\geq I(x^{\prime},y^{\prime})$. The orbital and pair entropies and consequently the mutual information (11,29) are computed from the eigenvalues of the density matrices, which fulfill all requirements on being interpreted as probability distributions, so the classical information theory results hold for these quantities. Since the spin summation (20) is a function mapping between two such distributions, the data processing inequality must hold. Intuitively, it is clear that if we discard some information by performing the spin summation, we cannot increase the amount of information correlating one orbital with another. The transition from the original mutual information (11) to the spin-free one (29) can thus only simplify the diagrams for the correlation analysis.
Another advantage of the spin-free quantities is their invariance with respect to the spin component $M_{s}$ of a given spin multiplet state. To show this, we assume that the same molecular orbitals have been employed in calculations of all the $M_{s}$ components of the multiplet and start from the well known fact that the total $N$-electron wave function can be written as a product of spatial and spin wave functions
@@
where $\Psi^{\prime}_{{\cal E}_{i}}(N-1)$ and $\Psi^{\prime\prime}_{{\cal E}_{i}}(N-1)$ are not in general spin eigenfunctions, the spin-trace over ${\cal E}_{i}$ does not form a full spin trace and the $(\rho_{i})_{22}$ and $(\rho_{i})_{33}$ matrix elements are not invariant individually. However, when taking their sum $(\rho_{i})_{22}+(\rho_{i})_{33}$, we sum over the spin variable of the electron in orbital $i$ as well, forming a full trace over all the spin variables
@@
which manifestly expresses the invariance of this quantity (and consequently of the spin-free entropy $\tilde{S}_{i}$) with respect to $M_{s}$. Actually, this equation might serve as an alternative definition of the spin-free orbital RDM, which would be particularly suitable in the context of spin-adapted methods like GUGA-CI. For the pair reduced density matrix, the $M_{s}$ invariance can be shown analogously, analyzing the 16 cases corresponding to the pair basis ${\cal B}\otimes{\cal B}$.
Notice, however, that neither the spin-free nor the spin-dependent orbital entropies and mutual informations are invariant with respect to orbital transformations, so the correlation pattern depends on the form of the orbitals (e.g. localization) [45, 41, 46] and a sensible (chemically intuitive) choice can be very important for its interpretation [47]. This non-invariance can actually be turned into an advantage, simplifying the multireference character of the wave function by a suitable orbital transformation. Of course, the orbital correlation cannot be fully transformed away by the change of on-particle basis, there exists an orbital basis for which it has a minimum value called particle correlation [41]. This idea dates back to Löwdin [48], who has shown that natural orbitals give rise to most compact CI expansion, but the relationship to quantum information theory was uncovered only recently. Legeza et al. have shown that by performing suitable orbital transformations (called fermionic mode optimization in DMRG [49]), it is possible to minimize the multireference character (i.e. the orbital correlation) in the wave function [50]. The orbital transformations are not limited to DMRG, but can be exploited in other strongly correlated methods as well. Li Manni et al. [51, 52, 53, 54] introduced a “Quantum Anamorphosis” approach in the spin-adapted GUGA-FCIQMC context and were able to achieve a dramatic compression of the multireference wave function by localization of the orbitals carrying the unpaired spins. They applied it to studies of polynuclear transition metal clusters, including the iron-sulfur compounds [36, 37], and by MO localization and reordering were able to achieve remarkable sparsity of the FCIQMC Hamiltonian matrices leading to improvements in computational efficiency as well as insight into the spin structure of the investigated states. Along similar lines, Schilling et al. [55] have employed orbital transformations minimizing the total orbital correlation to improve the ability of the tailored CCSD method to capture the correlation energy.
@@ 3 Computational details
For the illustration of the spin-free correlation analysis we employed a model example of a non-interacting dimer of CH2 diradicals in triplet ground states coupled to different resulting multiplicities. We took arbitrarily an equilateral triangular geometry with bond lengths 1.121Å  and angle 152.7 and computed FCI for the monomer (triplet ground state) employing canonical RHF orbitals in the STO-3G basis and freezing the 1s orbitals on carbon atoms. The dimer was constructed from two monomers separated by 100 a.u. distance, and through different choices of initial guess we converged the FCI to various linear combinations of the degenerate ground states with different spin multiplicities. For these small model systems we employed our in-house FCI (traditional string-based algorithm) and entropy analysis C++ code. The molecular orbitals were visualized using CHARMOL [56].
@@
As a more realistic application we have chosen the iron complexes [Fe(SCH3)4]- and [Fe2S2(SCH3)4]2- (cf. Fig. 1), which were recently employed by Neese et al. to demonstrate the new restricted open shell SCF code in Orca [35]. The geometries of the complexes were originally reported in Refs. [33, 34], but are available in the Supplementary Information of Ref. [35]. We employed the Orca program [57] and the Def2-TZVP basis set to perform a configuration-averaged ROHF SCF calculations and integral dump, while the MOLMPS code[58] was used for the subsequent DMRG calculations, and the new C++ code for the entropy analysis.
For the single-iron complex, the ROHF SCF was done configuration-averaged for 5 electrons in the 5 d-orbitals, yielding one orbital set, and separate ROHF calculations have been performed also for the three spin states. DMRG(13,14) calculations with bond dimension 512 were performed for states with multiplicities 2,4,6 in each of the four orbital sets. For the double-iron-sulfur complex, the configuration-averaged SCF was performed for the 10 electrons in iron d-orbitals and subsequently DMRG (14,14) and (22,21) calculations with bond dimension 2048 have been computed for the spin multiplicities 1,3,…,11. It turns out that in this complex the dynamic correlation is critical to get the correct energetic ordering of the states; indeed, in the (14,14) the high-spin state is incorrectly the lowest, while already in the (22,21) space the energy ordering was qualitatively correct (singlet ground state), in agreement with the post-DMRG and NEVPT2 calculations [29]. We did not perform computations for several values of bond dimension and extrapolation to infinite bond dimension, as we did not aim to compute the correlation energy with high accuracy, however, the active spaces and bond dimensions employed give a qualitatively correct energy ordering of the states and are sufficient for the qualitative correlation analysis. The DMRG calculations were not spin adapted, but we checked the spin contamination and the $S^{2}$ eigenvalues deviated from the correct ones at most by 0.002 for all calculations of the single-iron complex and at most by 0.06 for the lowest root of each $M_{s}$ in the two-iron complex, and at most 0.2 for its higher roots (while the eigenvalues themselves are up to 30). However, it was not possible to converge all excited states corresponding to low $M_{s}$ components of the high $S$ states.